OpenStudy (anonymous):

can you help on this one? 3/-4+4/9 i find the common denominator first right?

OpenStudy (anonymous):

Yes find the LCD Then solve by adding and simplifying.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

36?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so multiply 3/-4 by 9 right? to get 27/-36?

OpenStudy (anonymous):

Find the LCD (36) \[\frac{ -3 }{ 4 }+\frac{ 4 }{ 9 }\] Now multiply each term\[\frac{ -27 }{ 36 } + \frac{ 16 }{ 36 }\] Now we add. \[\frac{ -11 }{ 36 }\]